# Provably Secure Covert Messaging Using Image-based Diffusion Processes

### v1

@@ Abstract
Index Terms:
steganography, covert messaging, diffusion
@@ I Introduction
We focus on diffusion models [1, 2, 3, 4], which have recently emerged as state-of-the-art generative models for computer vision, due to reportedly outperforming generative adversarial networks (GANs) [5] in output quality [6]. Like GANs, we can view diffusion models as mapping a noise distribution into a desired distribution (e.g., images). However, sampling in diffusion models works by iterative denoising. Starting from a pure noise sample, the model transforms it into a realistic output through consecutive denoising steps.
@@
Our proposed construction has two other desirable features. First, it incorporates a novel error correction technique. This error correction mechanism is built on a specific inversion technique called EDICT [7], which performs inversion using dual latent spaces. We show how to leverage the dual latent spaces to obtain improved message recovery rates. Second, we show that embedding covert bits directly in the initial latents noise provides superior robustness to output image transformations (e.g., resize, blur, etc.) than alternatives.
Our paper’s contribution is perhaps best conceptualized in relation to the growing literature consisting of numerous prior attempts at doing steganography and (or) covert messaging with generative models such as with text/LLM [8, 9, 10, 11], speech [12, 13, 14], images GANs [15, 16, 17, 18, 19, 20, 21]. However, existing proposals using diffusion models [22, 23, 24, 15] lack robustness, require costly model retraining, and (or) are simply insecure.
@@
To demonstrate this, we focus on related work Kim et al. [15], the closest prior attempt to ours in terms of embedding strategy. They propose an embedding strategy they call message projection. They achieve the high output quality they strive for, but distort the latent noise distribution. We construct an attack that breaks their scheme, despite them achieving the standard steganographic ideal of output indistinguishability.
@@
  • Error correction: we devise an error correction scheme using EDICT [7]. This improves embedding capacity and increase robustness over the traditional DDIM scheduler.
@@ II-A Diffusion models
Diffusion models [1, 2, 3, 4] can be described through their forward and reverse processes. In the forward process, a clean image ${{\bm{x}}_{0}}$ is progressively corrupted by adding Gaussian noise through a series of $T$ steps from ${{\bm{x}}_{0}}$ to ${{\bm{x}}_{T}}$. The reverse process aims to reconstruct the original image ${{\bm{x}}_{0}}$ by iteratively denoising the noisy image ${{\bm{x}}_{T}}$. Starting from ${{\bm{x}}_{T}}$, the model denoises the image step-by-step, reducing the noise and eventually reconstructing the image at ${{\bm{x}}_{0}}$.
Forward & Reverse Processes. Among the various sampling techniques used in diffusion models, such as Denoising Diffusion Probabilistic Models (DDPM) [3] and Denoising Diffusion Implicit Models (DDIM) [4], deterministic samplers like DDIM are of particular interest. Deterministic samplers are preferred because they are invertible, allowing the reverse process to be precisely defined and controlled. The invertibility of DDIM is crucial for ensuring that the denoising steps can be accurately traced back, facilitating high-fidelity image generation. The formula of the forward process in DDIM is:
@@
In addition to DDIM, we use EDICT [7]. EDICT is a diffusion scheduler that follows a similar structure to DDIM, but creates two latent spaces. These are initially the same space but diverge during the denoising process as they are continually mixed with each other. This uses a similar estimation process that uses mixing to reduce inversion error, especially for models that use complex prompts.
Latent Diffusion Models (LDM). StableDiffusion [^2] https://stability.ai/ & https://huggingface.co/stabilityai/stable-diffusion-2-1 [25], an advanced application of diffusion models, operates as a Latent Diffusion Model (LDM). LDMs leverage a latent space representation that the diffusion process operates over. The result is that it enhances the model’s ability to generate detailed and high-resolution images. The key components of StableDiffusion include not only the diffusion process over a compressed 64x64 latent space but also an encoder and decoder to convert the latent representations back and forth between the full-size images, typically at resolutions such as 512x512. The encoder compresses the high-resolution images into the latent space, which the diffusion process then modifies. Following this, the decoder reconstructs the modified latent image back into a high-resolution output. The stability and efficiency of LDMs make them a preferred choice for complex image-generation tasks.
@@
In the sampling process, we start with a text prompt $t$ and an initial latent noise sample ${{\bm{x}}_{T}}$ (isotropic standard Gaussian [3]), which is represented as a 64x64x4 tensor where the last dimension is the number of channels (RGBA). The StableDiffusion sampling process will then iteratively denoise ${{\bm{x}}_{T}}$ into ${{\bm{x}}_{0}}$ (same dimension), and then a full resolution 512x512 image will be decoded from ${{\bm{x}}_{0}}$. Note that this process is completely deterministic once ${{\bm{x}}_{T}}$ and the prompt are determined. Given a full size resolution 512x512 image (which may or may not have been created through StableDiffusion) the inversion process first uses the encoder to obtain $\tilde{{{\bm{x}}_{0}}}$ a 64x64x4 tensor representation of the image. It then iteratively noises this representation back into a latent point $\tilde{{{\bm{x}}_{T}}}$. As mentioned before since the inversion process is only approximate, the inversion of a sampled initial latent will have lost information. In other words, ${{\bm{x}}_{T}}\neq\tilde{{{\bm{x}}_{T}}}$ although we expect ${{\bm{x}}_{T}}\approx\tilde{{{\bm{x}}_{T}}}$.
@@ II-B Generative Steganography
Since the advent of deep learning there has been significant interest in doing steganography using deep generative models such as LLMs for text-based steganography [11, 9, 8, 10], GANs for images [26, 27, 28], and audio/speech [29, 30]. In this method, the receiver extracts the message from the generated “cover media” through an inversion process or a separate model.
Several researchers have explored image steganography using neural network frameworks such as GAN [16] or Glow [17]. But were limited by image quality constraints.
Diffusion models have surpassed these other models. Jois et al. [31] propose a scheme called “Pulsar” for embedding secret data in diffusion model outputs, however, it is designed for pixel space diffusion models, not the more popular latent space models.
There are several schemes using latent space diffusion models, but their approaches vary. Peng et al. [22] developed “StegaDDPM”, embedding secret messages via denoising diffusion models. But this method suffers from minor distortions leading to inversion failures. Yu et al. [23] propose a method called “CRoSS”, a steganography framework using DDIM to convert cover to secret images, differing from traditional embedding methods. Wei et al. [24] propose Generative Steganography Diffusion, using an invertible model and DCT for data embedding. However, their approach requires significant resources to retrain the model. Peng et al. [32] proposed “LDStega,” a robust steganography technique for embedding confidential data within images. They demonstrate its resilience to various image distortions, including Gaussian noise and compression.
In contrast to existing techniques, our method uses off-the-shelf latent diffusion models such as StableDiffusion without the need for retraining. It also embeds covert messages in the initial latents (“latent noise”) rather than later on in the diffusion process to obtain greater robustness to modification of the output (e.g., resize, compression, etc.).
Given these desiderata, the work most closely related to ours is Kim et al. [15]. They introduce three distinct message projection methods, each designed to embed secret messages into latent noise. The major limitation of their technique is that it alters the latent space distribution. This may not seem to be a significant issue as their outputs appear realistic. But in fact (as we show empirically) it leads to a total break. An adversary can efficiently detect their cover messages with probability almost 1. By contrast, our proposed technique exactly preserves the latent space distribution and we prove it achieves indistinguishability.
@@
Peng et al.[22] $\times$ $\times$ $\times$ $\times$
Yu et al.[23] $\times$ $\times$ $\times$ $\times$
Wei et al.[24] $\times$ $\times$ $\times$ $\times$ $\times$
Peng et al. [32] $\times$ $\times$ $\times$ $\times$
Kim et al. [15] $\times$ $\times$ $\times$ $\times$
@@ II-C Watermarking
A different problem that relies on similar techniques to ours is watermarking; in which a model owner embeds a small watermark within the generated diffusion output. This problem has received substantial attention recently [33, 34, 35, 36]. In particular, two methods also attempt to invisibly embed information in the latent space. Tree Ring Watermarking [33] embeds in patterns structured in Fourier space that are highly robust, but influence model output. Gaussian Shading [34] seeks to match the natural Gaussian distribution, and achieves high robustness, but has limited capacity. We discuss these papers because watermarking has similarities to generative steganography in that both attempt to embed information into generated images without decreasing the quality of the output. However, they differ in their goals and threat models. Watermarking strives for easy retrieval and robustness against alterations. By contrast, the goal of steganography is to create a reliable channel between the sender and receiver that is completely undetectable and maximizes its capacity.
@@ IV-A Initial Setup
2:${{\bm{x}}_{T}}^{\prime}$: Latents with embedded ciphertext $c$
3:$c^{\prime}\leftarrow{\rm AddRedundancy}(c,\rho)$triangleright Within $c$ convert all 0 bits to -1, then apply redundancy by repeating each bit $\rho$ times.
4:${\rm c_{m}^{\prime}}\leftarrow{\rm Permute}(c^{\prime},\pi)$triangleright Permute $c^{\prime}$, according to $\pi$
5:${{\bm{x}}_{T}}^{\prime}\leftarrow{{\bm{x}}_{T}}$
@@
9:         if $s=-1$ then
10:             ${{x}_{T,{i}}}\leftarrow-{{x}_{T,{i}}}$triangleright Flip the sign of $i^{\rm th}$ element of ${{\bm{x}}_{T}}^{\prime}$
@@ IV-C Ciphertext Embedding
Recall that due to their shared key material, Alice and Bob have agreed on parameters. For StableDiffusion, the initial latents ${{\bm{x}}_{T}}$ are represented as 64x64x4 tensors sampled as isotropic Gaussian [3]. Said differently, each latent component ${{x}_{T,{i}}}$ is sampled i.i.d. from $\mathcal{N}(0,1)$. So the only difference in this case is that the PRNG is seeded based on key material derived from the shared secret key and the message counter. This way Alice and Bob will both obtain the same initial latents ${{\bm{x}}_{T}}$ for each message and can use a similar process to generate the same permutation $\pi$.
Given this, Alice has to perform two steps for embedding a ciphertext $c=c_{\rm body}||c_{\rm tag}$. First, redundancy is added to the ciphertext as follows. Each bit in $c$ is repeated $\rho\geq 1$ times (redundancy parameter) before permuting the (redundant) ciphertext according to permutation $\pi$, creating $c^{\prime}_{m}$. Second, we only embed bits into component values (of ${{\bm{x}}_{T}}$) with absolute values above a certain threshold $\tau\geq 0$. So for each component value within the 64x64x4 latent space, if its absolute value exceeds $\tau$ we embed a bit from $c^{\prime}_{m}$. To embed a $1$ bit, we do not change the component value, and to embed a 00 bit, we flip the current sign of the component value. As we show in Section V-D, this flipping of bits does not alter the distribution. The method is illustrated in Algorithm 1. Lines 7-8 perform the actual logic embedding bits. The process can be thought of as applying a mask of permuted values over valid values. If we have redundancy $\rho$ then we can only embed $16384/\rho$ bits of the ciphertext at the most per output (assuming 64x64x4 latent space), and any component value that does not exceed $\tau$ in magnitude is not embedded with a bit. We discuss this further in Section VIII.
@@
1:${{\bm{x}}_{T}}$: Initial latents; $\pi$: Permutation; $\tilde{{{\bm{x}}_{T}}}$: Reconstructed latents; $\rho$: Redundancy parameter; $\tau$: Threshold parameter; $p(\cdot)$: pdf of $\mathcal{N}(0,1)$; $\triangleright$ Assume errors are standard normal
2:$\tilde{c}$: reconstructed ciphertext
@@
10:    $\rm v^{\prime}\leftarrow{\rm Inverse\_permute}(\rm v,\pi,\tau)$triangleright Given a list of latent space values remove the applied Mask by identifying all values with a magnitude greater than $\tau$, and then inverting the permutation. Returning latent space values respective to $c^{\prime}$ ($c$ with redundancy applied)
11:    $\rm v^{\prime}\leftarrow{\rm Partition}(v,\rho)$triangleright Partition list into a list of $\rho$ length lists such that sub-list $i$ contains all the latent space values embedded with the $i$th bit of $c$
@@
14:${\rm v_{\rm orig}}\leftarrow{\rm Flatten}({{\bm{x}}_{T}})$
15:${\rm v_{\rm recov}}\leftarrow{\rm Flatten}(\tilde{}{{\bm{x}}_{T}})$
16:${\rm v_{orig}^{\prime}}\leftarrow{\rm Remove\_Mask}({\rm v_{orig}},\pi,\rho,\tau)$
17:${\rm v_{recov}^{\prime}}\leftarrow{\rm Remove\_Mask}({\rm v_{recov}},\pi,\rho,\tau)$
18:$\tilde{c}\leftarrow``"$
@@ IV-D Latent Space Inversion
Our construction uses the DDIM scheduler inversion process proposed by Ho et al. [3] and as such can be used with any diffusion model using that scheduler. In this paper, in addition to using the DDIM scheduler, we also propose to use EDICT, a variant of DDIM that uses dual latent spaces to improve inversion accuracy [7]. More specifically, when denoising at each step, the scheduler will update one latent space with information from the other one, in a deterministic and invertible way. This helps stabilize the inversion process, allowing us to perform error correction. The use of EDICT increases message accuracy (the percentage of messages sent that can be perfectly recovered) as we detail in Section VIII. When embedding, the only difference between EDICT and DDIM is that we duplicate ${{\bm{x}}_{T}}^{\prime}$, to form $({{\bm{x}}_{T}}^{\prime},{{\bm{x}}_{T}}^{\prime})$ which functions as the dual latent spaces that initiate image generation. On the receiving end we perform the same retrieval process across both latent spaces, and then use them to perform error correction as detailed in Algorithm 3.

### v2

@@ I Introduction
We focus on diffusion models [sohl2015deep, song2019generative, ho2020denoising, song2020denoising], which have recently emerged as state-of-the-art generative models for computer vision, due to reportedly outperforming generative adversarial networks (GANs) [goodfellow2020generative] in output quality [dhariwal2021diffusion]. Like GANs, we can view diffusion models as mapping a noise distribution into a desired distribution (e.g., images). However, sampling in diffusion models works by iterative denoising. Starting from a pure noise sample, the model transforms it into a realistic output through consecutive denoising steps.
@@
Our proposed construction has two other desirable features. First, it incorporates a novel error correction technique. This error correction mechanism is built on a specific inversion technique called EDICT [wallace2023edict], which performs inversion using dual latent spaces. We show how to leverage the dual latent spaces to obtain improved message recovery rates. Second, we show that embedding covert bits directly in the initial latents noise provides superior robustness to output image transformations (e.g., resize, blur, etc.) than alternatives.
Our paper’s contribution is perhaps best conceptualized in relation to the growing literature consisting of numerous prior attempts at doing steganography and (or) covert messaging with generative models such as with text/LLM [bauer2021covert, kaptchuk2021meteor, cao2022generative, dai2019towards], speech [dutta2020overview, li2023coverless, chen2021distribution], images GANs [kim2023diffusion, wei2022generative, zhou2022secret, zhu2017unpaired, qin2019coverless, liu2017coverless, zhang2020generative]. However, existing proposals using diffusion models [peng2023stegaddpm, yu2024cross, wei2023generative, kim2023diffusion] lack robustness, require costly model retraining, and (or) are simply insecure.
@@
To demonstrate this, we focus on related work Kim et al. [kim2023diffusion], the closest prior attempt to ours in terms of embedding strategy. They propose an embedding strategy they call message projection. They achieve the high output quality they strive for, but distort the latent noise distribution. We construct an attack that breaks their scheme, despite them achieving the standard steganographic ideal of output indistinguishability.
@@
  • Error correction: we devise an error correction scheme using EDICT [wallace2023edict]. This improves embedding capacity and increase robustness over the traditional DDIM scheduler.
@@ II-A Diffusion models
Diffusion models [sohl2015deep, song2019generative, ho2020denoising, song2020denoising] can be described through their forward and reverse processes. In the forward process, a clean image ${\bm{x}}_{0}$} is progressively corrupted by adding Gaussian noise through a series of $T$ steps from ${\bm{x}}_{0}$} to ${\bm{x}}_{T}$}. The reverse process aims to reconstruct the original image ${\bm{x}}_{0}$} by iteratively denoising the noisy image ${\bm{x}}_{T}$}. Starting from ${\bm{x}}_{T}$}, the model denoises the image step-by-step, reducing the noise and eventually reconstructing the image at ${\bm{x}}_{0}$}.
Forward & Reverse Processes. Among the various sampling techniques used in diffusion models, such as Denoising Diffusion Probabilistic Models (DDPM) [ho2020denoising] and Denoising Diffusion Implicit Models (DDIM) [song2020denoising], deterministic samplers like DDIM are of particular interest. Deterministic samplers are preferred because they are invertible, allowing the reverse process to be precisely defined and controlled. The invertibility of DDIM is crucial for ensuring that the denoising steps can be accurately traced back, facilitating high-fidelity image generation. The formula of the forward process in DDIM is:
@@
In addition to DDIM, we use EDICT [wallace2023edict]. EDICT is a diffusion scheduler that follows a similar structure to DDIM, but creates two latent spaces. These are initially the same space but diverge during the denoising process as they are continually mixed with each other. This uses a similar estimation process that uses mixing to reduce inversion error, especially for models that use complex prompts.
Latent Diffusion Models (LDM). StableDiffusion [^2] https://stability.ai/ & https://huggingface.co/stabilityai/stable-diffusion-2-1 [rombach2022high], an advanced application of diffusion models, operates as a Latent Diffusion Model (LDM). LDMs leverage a latent space representation that the diffusion process operates over. The result is that it enhances the model’s ability to generate detailed and high-resolution images. The key components of StableDiffusion include not only the diffusion process over a compressed 64x64 latent space but also an encoder and decoder to convert the latent representations back and forth between the full-size images, typically at resolutions such as 512x512. The encoder compresses the high-resolution images into the latent space, which the diffusion process then modifies. Following this, the decoder reconstructs the modified latent image back into a high-resolution output. The stability and efficiency of LDMs make them a preferred choice for complex image-generation tasks.
@@
In the sampling process, we start with a text prompt $t$ and an initial latent noise sample ${\bm{x}}_{T}$} (isotropic standard Gaussian [ho2020denoising]), which is represented as a 64x64x4 tensor where the last dimension is the number of channels (RGBA). The StableDiffusion sampling process will then iteratively denoise ${\bm{x}}_{T}$} into ${\bm{x}}_{0}$} (same dimension), and then a full resolution 512x512 image will be decoded from ${\bm{x}}_{0}$}. Note that this process is completely deterministic once ${\bm{x}}_{T}$} and the prompt are determined. Given a full size resolution 512x512 image (which may or may not have been created through StableDiffusion) the inversion process first uses the encoder to obtain $\tilde{{{\bm{x}}_{0}}}$ a 64x64x4 tensor representation of the image. It then iteratively noises this representation back into a latent point $\tilde{{{\bm{x}}_{T}}}$. As mentioned before since the inversion process is only approximate, the inversion of a sampled initial latent will have lost information. In other words, ${{\bm{x}}_{T}}\neq\tilde{{{\bm{x}}_{T}}}$ although we expect ${{\bm{x}}_{T}}\approx\tilde{{{\bm{x}}_{T}}}$.
@@ II-B Generative Steganography
Since the advent of deep learning there has been significant interest in doing steganography using deep generative models such as LLMs for text-based steganography [dai2019towards, kaptchuk2021meteor, bauer2021covert, cao2022generative], GANs for images [zhang2019steganogan, hu2018novel, liu2020recent], and audio/speech [yang2020approaching, ye2019heard]. In this method, the receiver extracts the message from the generated “cover media” through an inversion process or a separate model.
Several researchers have explored image steganography using neural network frameworks such as GAN [wei2022generative] or Glow [zhou2022secret]. But were limited by image quality constraints.
Diffusion models have surpassed these other models. Jois et al. [jois2023pulsar] propose a scheme called “Pulsar” for embedding secret data in diffusion model outputs, however, it is designed for pixel space diffusion models, not the more popular latent space models.
There are several schemes using latent space diffusion models, but their approaches vary. Peng et al. [peng2023stegaddpm] developed “StegaDDPM”, embedding secret messages via denoising diffusion models. But this method suffers from minor distortions leading to inversion failures. Yu et al. [yu2024cross] propose a method called “CRoSS”, a steganography framework using DDIM to convert cover to secret images, differing from traditional embedding methods. Wei et al. [wei2023generative] propose Generative Steganography Diffusion, using an invertible model and DCT for data embedding. However, their approach requires significant resources to retrain the model. Peng et al. [peng2024ldstega] proposed “LDStega,” a robust steganography technique for embedding confidential data within images. They demonstrate its resilience to various image distortions, including Gaussian noise and compression. Concurrent work by Mahfuz et al. [mahfuz2025psyduck] is PSyDUCK. This method embeds a message by diverging the generation process at certain timesteps in the denoising process; it then mixes the diverged samples back together using the covert message. Observing the final image, the receiver can identify the message used to mix the divergent paths. They apply this method to pixel-based and latent space diffusion models for images and videos.
In contrast to existing techniques, our method both uses off-the-shelf latent diffusion models without retraining, and it embeds covert messages in the initial latents (“latent noise”) rather than later on in the diffusion process to obtain greater robustness to modification of the output (e.g., resize, compression, etc.).
Given these desiderata, the work most closely related to ours is Kim et al. [kim2023diffusion]. They introduce three distinct message projection methods, each designed to embed secret messages into latent noise. The major limitation of their technique is that it alters the latent space distribution. This may not seem to be a significant issue as their outputs appear realistic. But in fact (as we show empirically) it leads to a total break. An adversary can efficiently detect their cover messages with probability almost 1. By contrast, our proposed technique exactly preserves the latent space distribution and we prove it achieves indistinguishability.
@@
Peng et al.[peng2023stegaddpm] $\times$ $\times$ $\times$ $\times$
Yu et al.[yu2024cross] $\times$ $\times$ $\times$ $\times$
Wei et al.[wei2023generative] $\times$ $\times$ $\times$ $\times$ $\times$
Peng et al. [peng2024ldstega] $\times$ $\times$ $\times$ $\times$
Kim et al. [kim2023diffusion] $\times$ $\times$ $\times$ $\times$
@@ II-C Watermarking
A different problem that relies on similar techniques to ours is watermarking; in which a model owner embeds a small watermark within the generated diffusion output. This problem has received substantial attention recently [wen2023tree, yang2024gaussian, liu2023watermarking, zhao2023recipe]. In particular, two methods also attempt to invisibly embed information in the latent space. Tree Ring Watermarking [wen2023tree] embeds in patterns structured in Fourier space that are highly robust, but influence model output. Gaussian Shading [yang2024gaussian] seeks to match the natural Gaussian distribution, and achieves high robustness, but has limited capacity. We discuss these papers because watermarking has similarities to generative steganography in that both attempt to embed information into generated images without decreasing the quality of the output. However, they differ in their goals and threat models. Watermarking strives for easy retrieval and robustness against alterations. By contrast, the goal of steganography is to create a reliable channel between the sender and receiver that is completely undetectable and maximizes its capacity.
@@ IV-A Initial Setup
[…]
@@
9:    if s=1s=-1 then
10:     ${{x}_{T,{i}}}\leftarrow-{{x}_{T,{i}}}$ $\triangleright$ Flip the sign of $i^{\rm th}$ element of ${\bm{x}}_{T}$$^{\prime}$
@@ IV-C Ciphertext Embedding
Recall that due to their shared key material, Alice and Bob have agreed on parameters. For StableDiffusion, the initial latents ${\bm{x}}_{T}$} are represented as 64x64x4 tensors sampled as isotropic Gaussian [ho2020denoising]. Said differently, each latent component ${x}_{T,{i}}$} is sampled i.i.d. from $\mathcal{N}(0,1)$. So the only difference in this case is that the PRNG is seeded based on key material derived from the shared secret key and the message counter. This way Alice and Bob will both obtain the same initial latents ${\bm{x}}_{T}$} for each message and can use a similar process to generate the same permutation $\pi$.
Given this, Alice has to perform two steps for embedding a ciphertext $c=c_{\rm body}||c_{\rm tag}$. First, redundancy is added to the ciphertext as follows. Each bit in $c$ is repeated $\rho\geq 1$ times (redundancy parameter) before permuting the (redundant) ciphertext according to permutation $\pi$, creating $c^{\prime}_{m}$. Second, we only embed bits into component values (of ${\bm{x}}_{T}$}) with absolute values above a certain threshold $\tau\geq 0$. So for each component value within the 64x64x4 latent space, if its absolute value exceeds $\tau$ we embed a bit from $c^{\prime}_{m}$. To embed a $1$ bit, we do not change the component value, and to embed a 0 bit, we flip the current sign of the component value. As we show in Section V-D, this flipping of bits does not alter the distribution. The method is illustrated in Algorithm 1. Lines 7-8 perform the actual logic embedding bits. The process can be thought of as applying a mask of permuted values over valid values. If we have redundancy $\rho$ then we can only embed $16384/\rho$ bits of the ciphertext at the most per output (assuming 64x64x4 latent space), and any component value that does not exceed $\tau$ in magnitude is not embedded with a bit. We discuss this further in Section VIII.
@@
[…]
@@
10:  $\rm v^{\prime}\leftarrow{\rm Inverse\_permute}(\rm v,\pi,\tau)$ $\triangleright$ Given a list of latent space values remove the applied Mask by identifying all values with a magnitude greater than $\tau$, and then inverting the permutation. Returning latent space values respective to $c^{\prime}$ ($c$ with redundancy applied)
11:  $\rm v^{\prime}\leftarrow{\rm Partition}(v,\rho)$ $\triangleright$ Partition list into a list of $\rho$ length lists such that sub-list $i$ contains all the latent space values embedded with the $i$th bit of $c$
@@
[…]
@@ IV-D Latent Space Inversion
Our construction uses the DDIM scheduler inversion process proposed by Ho et al. [ho2020denoising] and as such can be used with any diffusion model using that scheduler. In this paper, in addition to using the DDIM scheduler, we also propose to use EDICT, a variant of DDIM that uses dual latent spaces to improve inversion accuracy [wallace2023edict]. More specifically, when denoising at each step, the scheduler will update one latent space with information from the other one, in a deterministic and invertible way. This helps stabilize the inversion process, allowing us to perform error correction. The use of EDICT increases message accuracy (the percentage of messages sent that can be perfectly recovered) as we detail in Section VIII. When embedding, the only difference between EDICT and DDIM is that we duplicate ${\bm{x}}_{T}$$^{\prime}$, to form $({{\bm{x}}_{T}}^{\prime},{{\bm{x}}_{T}}^{\prime})$ which functions as the dual latent spaces that initiate image generation. On the receiving end we perform the same retrieval process across both latent spaces, and then use them to perform error correction as detailed in Algorithm 3.